# Normalized Remainders: Definition, Origins, Backgrounds, Recent Results, and Basic Properties

### v1

@@ 3.3. Normalized remainders of exponential function
for $n\in\mathbb{N}_{0}$ and $x\in\mathbb{R}$, and making use of two monotonicity rules in [2, pp. 10–11, Theorem 1.25] and in [48, Lemma 9], [53, Lemma 12], and [58, Remark 7.2], respectively, we concisely and elegantly proved the following properties of the normalized remainder $T_{n}[\operatorname{e}^{x}]$ once again:
  1. (1)
    For $n\in\mathbb{N}_{0}$, the normalized remainder $T_{n}[\operatorname{e}^{x}]$ is positive in $x\in\mathbb{R}$. See also [47, Section 2].
  2. (2)
    For $n\in\mathbb{N}_{0}$, the normalized remainder $T_{n}[\operatorname{e}^{x}]$ is an increasing and logarithmically convex function of $x\in\mathbb{R}$. See also [5, Corollary 1].
  3. (3)
    For $n\in\mathbb{N}_{0}$, the normalized remainder $T_{n}[\operatorname{e}^{x}]$ is an absolutely monotonic function in $x\in\mathbb{R}$. See also [47, Theorem 2].
  4. (4)
    For $n\in\mathbb{N}$, the function $R_{n,0}(x)$ defined by (3.10) is increasing in $x\in\mathbb{R}$. See also [5, Theorem 2].
  5. (5)
    For $x\in\mathbb{R}$ and $n\in\mathbb{N}_{0}$, the inequalities (3.11) and (3.12) with equality are true. See also [5, Corollary 2] and [78, Theorem 1].
  6. (6)
    For $n\in\mathbb{N}_{0}$, the ratio $\frac{T_{n+1}[\operatorname{e}^{x}]}{T_{n}[\operatorname{e}^{x}]}$ is decreasing in $x\in\mathbb{R}$. See also [47, Theorem 1] and [78, Theorem 2].
@@ 3.4. Normalized remainders of inverse trigonometric functions
$T_{2n}[\arcsin x]=\begin{dcases}\frac{[(2n)!!]^{2}}{(2n)!}\frac{2n+1}{x^{2n+1}}\Biggl(\arcsin x-\sum_{k=0}^{n-1}\frac{(2k)!}{[(2k)!!]^{2}}\frac{x^{2k+1}}{2k+1}\Biggr),&x\neq 0\\ 1,&x=0\end{dcases}$
@@
$T_{2n}[\arctan x]=\begin{dcases}(-1)^{n}\frac{2n+1}{x^{2n+1}}\Biggl[\arctan x-\sum_{k=0}^{n-1}(-1)^{k}\frac{x^{2k+1}}{2k+1}\Biggr],&x\neq 0\\ 1,&x=0\end{dcases}$
for $n\in\mathbb{N}_{0}$. We derived the following results:
  1. (1)
    For $n\in\mathbb{N}_{0}$, the normalized remainders $T_{2n}[\arcsin x]=T_{2n}[\arccos x]$ and $T_{2n}[\arctan x]$ are positive and even in $x\in(-1,1)$.
  2. (2)
    For $n\in\mathbb{N}_{0}$, the normalized remainders $T_{2n}[\arcsin x]=T_{2n}[\arccos x]$ are increasing and the normalized remainder $T_{2n}[\arctan x]$ is decreasing in $x\in(0,1)$.
  3. (3)
    For $n\in\mathbb{N}_{0}$, the normalized remainder $T_{2n}[\arcsin x]$ is logarithmically convex in $x\in(0,1)$.
  4. (4)
    For $n\in\mathbb{N}_{0}$ and $|x|<1$, the normalized remainder
    $T_{2n}[\operatorname{arctanh}x]=\begin{dcases}\frac{2n+1}{x^{2n+1}}\Biggl(\operatorname{arctanh}x-\sum_{k=0}^{n-1}\frac{x^{2k+1}}{2k+1}\Biggr),&x\neq 0\\ 1,&x=0\end{dcases}$
    is logarithmically convex in $x\in(0,1)$.
  5. (5)
    The functions $\ln\bigl(\frac{2}{\pi}\arccos x\bigr)$, $\ln T_{2n}[\arctan x]$, and $\ln T_{2n}[\arcsin x]$ were expanded into the Maclaurin series.
  6. (6)
    For $n\in\mathbb{N}_{0}$, the ratio $\frac{T_{2n}[\arcsin x]}{T_{2n+2}[\arcsin x]}$ is decreasing in $x\in(0,1)$.
@@
In [68, p. 71, Eq. (3.38)], we find the integral representation
@@
where the classical Euler gamma function $\Gamma(z)$ can be defined [68, Chapter 3] by
@@
This can also be derived by the Laurent series expansion of $\cot\theta$ in the handbooks [1, p. 75, Entry 4.3.70] and [42, p. 117, Entry 4.19.6]. See also [79, Remark 1]. As a result, the function $1-\frac{\theta}{\tan\theta}$ is absolutely monotonic on $(0,\pi)$ and is completely monotonic on $(-\pi,0)$.
@@
  1. (1)
    The normalized remainder $T_{2n+1}[\Phi(\theta)]$ for $n\in\mathbb{N}_{0}$ is positive in $\theta\in(-\pi,\pi)$, is decreasing in $\theta\in(-\pi,0)$, and is increasing in $\theta\in(0,\pi)$.
  2. (2)
@@
  3. (3)
    The normalized remainder $T_{2n+1}[\Phi(\theta)]$ for $n\in\mathbb{N}_{0}$ is an absolutely monotonic function in $\theta\in(0,\pi)$ and is a completely monotonic function in $\theta\in(-\pi,0)$.
  4. (4)
@@
  5. (5)
    The ratio $\frac{T_{2n+3}[\Phi(\theta)]}{T_{2n+1}[\Phi(\theta)]}$ for $n\in\mathbb{N}_{0}$ is decreasing in $\theta\in(-\pi,0)$ and is increasing in $\theta\in(0,\pi)$.
@@
    1. (a)
      is an absolutely monotonic function in $\theta\in(0,\pi)$,
    2. (b)
      is a completely monotonic function in $\theta\in(-\pi,0)$,
    3. (c)
@@
The absolute and complete monotonicity of the ratio $\frac{T_{2n+3}[\Phi(\theta)]}{T_{2n+1}[\Phi(\theta)]}$ discovered in [65, Theorem 2] and the Laurent series of the function $\frac{1}{\Phi(\theta)}$ for $0<|\theta|<\pi$ expanded in [65, Theorem 3] imply that the function $\frac{2}{\theta^{2}}-\frac{1}{\Phi(\theta)}$ is absolutely monotonic on $(0,\pi)$ and completely monotonic on $(-\pi,0)$.
@@ 3.6. Normalized remainders of complete elliptic integrals
For given $x\in[0,1]$, the complete elliptic integral of the second kind $\mathscr{E}(x)$ are defined [68, p. 315] by
@@
In the paper [71, Remarks 2 and 4], we proved that the normalized remainders $T_{n}[\mathcal{H}(x)]$ and $T_{n}[\mathcal{G}(x)]$ for $n\in\mathbb{N}_{0}$ are absolutely monotonic on $(0,1)$.
@@
In the paper [23, Theorems 3 and 4], we employed a new method to concisely prove the absolute monotonicity of the normalized remainders $T_{n}[\mathcal{H}(x)]$ and $T_{n}[\mathcal{G}(x)]$ for $n\in\mathbb{N}_{0}$ on $(0,1)$ once again.
@@ 3.7. Normalized remainder of power of inverse sine function
In [17, Theorem 1], [18, Theorem 2.1], [49, Section 4], and [51, Section 6], we discovered the Maclaurin power series expansion
@@
for $n,\ell\in\mathbb{N}$ and the Stirling numbers of the first kind $s(\ell+n,n)$ for $\ell,n\in\mathbb{N}_{0}$ can be analytically generated [38, Theorem 3.14] by
@@
for $|x|<1$. These series expansions can be found in the literature [3, p. 444], [17, Remark 6], [18, Remark 5.2], [50, Section 1.1], [56, Section 1], [60, Section 1], [69, pp. 279 and 331–333], [82, p. 229], and [4, 8, 66], for example.
In [49, Theorem 7], see also [59, Example 3.1], the Maclaurin series expansion (3.15) was generalized as
@@
    where the central factorial numbers of the first kind $t(n,\ell)$ are defined [63, pp. 212–217, Section 6.5] by
@@ 4. Historic backgrounds in combinatorics and number theory
In this section, we collected some concepts and notions, which can be reformulated in terms of normalized remainders, in combinatorics and number theory.
@@ 4.1. Generating functions of Carlitz–Howard’s generalizations of Bernoulli numbers and polynomials
The generalized Bernoulli polynomials $B_{n}^{(\sigma)}(t)$ can be generated [68, p. 4] by
@@
for $n\in\mathbb{N}$, and they examined a lot of algebraic properties of these sequences and polynomials. It is easy to see that $A_{k}(0)=A_{k}$, $A_{1,k}=B_{k}$, $A_{2,k}=A_{k}$, $A_{n,k}(0)=A_{n,k}$, $A_{1,k}(0)=B_{k}$, $A_{2,k}(0)=A_{k}$, and $A_{1,k}(t)=B_{k}(t)$ for $k\in\mathbb{N}_{0}$. The sequence $A_{k}$ for $k\in\mathbb{N}_{0}$ was recently investigated in [5, Remark 2] and [61, Section 2]. In terms of the notation $T_{n-1}[\operatorname{e}^{x}]$, the equation (4.3) can be reformulated as
@@
See also [21, Remark 4], [22, Remark 4], [31, pp. 1036–1037], and [71, Remark 2].

### v2

@@ 3.3. Normalized remainders of exponential function
for $n\in\mathbb{N}_{0}$ and $x\in\mathbb{R}$, and making use of two monotonicity rules in [2, pp. 10–11, Theorem 1.25] and in [50, Lemma 9], [56, Lemma 12], and [61, Remark 7.2], respectively, we concisely and elegantly proved the following properties of the normalized remainder $T_{n}[\operatorname{e}^{x}]$ once again:
  1. 1)
    For $n\in\mathbb{N}_{0}$, the normalized remainder $T_{n}[\operatorname{e}^{x}]$ is positive in $x\in\mathbb{R}$; see also [49, Section 2].
  2. 2)
    For $n\in\mathbb{N}_{0}$, the normalized remainder $T_{n}[\operatorname{e}^{x}]$ is an increasing and logarithmically convex function of $x\in\mathbb{R}$; see also [5, Corollary 1].
  3. 3)
    For $n\in\mathbb{N}_{0}$, the normalized remainder $T_{n}[\operatorname{e}^{x}]$ is an absolutely monotonic function in $x\in\mathbb{R}$; see also [49, Theorem 2].
  4. 4)
    For $n\in\mathbb{N}$, the function $R_{n,0}(x)$ defined by (3.10) is increasing in $x\in\mathbb{R}$; see also [5, Theorem 2].
  5. 5)
    For $x\in\mathbb{R}$ and $n\in\mathbb{N}_{0}$, the inequalities (3.11) and (3.12) with equality are true; see also [5, Corollary 2] and [81, Theorem 1].
  6. 6)
    For $n\in\mathbb{N}_{0}$, the ratio $\frac{T_{n+1}[\operatorname{e}^{x}]}{T_{n}[\operatorname{e}^{x}]}$ is decreasing in $x\in\mathbb{R}$; see also [49, Theorem 1] and [81, Theorem 2].
3.3.6.
In the paper [23], by establishing the integral representations
$S_{m,n}(x)=\sum_{j=0}^{\infty}\frac{1}{\binom{j+m+n}{n}}\frac{x^{j}}{j!}=n\int_{0}^{1}v^{n-1}(1-v)^{m}\operatorname{e}^{x(1-v)}\operatorname{d\!}v$ (3.14)
and
$S_{\alpha,\beta}(x)=\sum_{j=0}^{\infty}\frac{1}{\binom{j+\alpha+\beta}{\beta}}\frac{x^{j}}{j!}=\beta\int_{0}^{1}v^{\beta-1}(1-v)^{\alpha}\operatorname{e}^{u(1-v)}\operatorname{d\!}v$
for $m,n\in\mathbb{N}_{0}$, $\Re(\alpha)>-1$, $\Re(\beta)>0$, and $x\in\mathbb{R}$, and by finding the relation
$S_{m,n+1}(x)=(T_{n}[\operatorname{e}^{x}])^{(m)}$
for $x\in\mathbb{R}$ and $m,n\in\mathbb{N}_{0}$, we reformulated the inequalities (3.11), (3.12), and
$\Biggl[\sum_{j=0}^{\infty}\frac{j+1}{(j+n+2)!}x^{j}\Biggr]\Biggl[\sum_{j=0}^{\infty}\frac{j+1}{(j+n)!}x^{j}\Biggr]<\Biggl[\sum_{j=0}^{\infty}\frac{j+1}{(j+n+1)!}x^{j}\Biggr]^{2}$
for $n\in\mathbb{N}$ and $x\in\mathbb{R}$ in [81, Lemma 3] as integral forms and in terms of $S_{m,n}(x)$, alternatively proved the absolute monotonicity in $x\in\mathbb{R}$ of the normalized remainder $T_{n}[\operatorname{e}^{x}]$ for $n\in\mathbb{N}_{0}$, and posed the following problems.
Problem 3.1 ([23, Problem 1]).
For $m,n\in\mathbb{N}$ and $x\in\mathbb{R}$, calculate the limits
$\lim_{n\to\infty}\frac{S_{m,n}(x)}{S_{m-1,n}(x)}=\lim_{n\to\infty}\frac{\int_{0}^{1}v^{n-1}(1-v)^{m}\operatorname{e}^{-xv}\operatorname{d\!}v}{\int_{0}^{1}v^{n-1}(1-v)^{m-1}\operatorname{e}^{-xv}\operatorname{d\!}v}$
and
$\lim_{m\to\infty}\frac{S_{m,n}(x)}{S_{m-1,n}(x)}=\lim_{m\to\infty}\frac{\int_{0}^{1}v^{n-1}(1-v)^{m}\operatorname{e}^{-xv}\operatorname{d\!}v}{\int_{0}^{1}v^{n-1}(1-v)^{m-1}\operatorname{e}^{-xv}\operatorname{d\!}v}.$
Problem 3.2 ([23, Problem 2]).
For $\Re(\alpha)>0$, $\Re(\beta)>0$, and $x\in\mathbb{R}$, compute the limits
$\lim_{\beta\to\infty}\frac{S_{\alpha,\beta}(x)}{S_{\alpha-1,\beta}(x)}=\lim_{\beta\to\infty}\frac{\int_{0}^{1}v^{\beta-1}(1-v)^{\alpha}\operatorname{e}^{-xv}\operatorname{d\!}v}{\int_{0}^{1}v^{\beta-1}(1-v)^{\alpha-1}\operatorname{e}^{-xv}\operatorname{d\!}v}$
and
$\lim_{\alpha\to\infty}\frac{S_{\alpha,\beta}(x)}{S_{\alpha-1,\beta}(x)}=\lim_{\alpha\to\infty}\frac{\int_{0}^{1}v^{\beta-1}(1-v)^{\alpha}\operatorname{e}^{-xv}\operatorname{d\!}v}{\int_{0}^{1}v^{\beta-1}(1-v)^{\alpha-1}\operatorname{e}^{-xv}\operatorname{d\!}v}.$
@@ 3.4. Normalized remainders of inverse trigonometric functions
$T_{2n}[\arcsin x]$ $=T_{2n}[\arccos x]$
$=\begin{dcases}\frac{[(2n)!!]^{2}}{(2n)!}\frac{2n+1}{x^{2n+1}}\Biggl(\arcsin x-\sum_{k=0}^{n-1}\frac{(2k)!}{[(2k)!!]^{2}}\frac{x^{2k+1}}{2k+1}\Biggr),&x\neq 0;\\ 1,&x=0,\end{dcases}$
$T_{2n}[\arctan x]$ $=\begin{dcases}(-1)^{n}\frac{2n+1}{x^{2n+1}}\Biggl[\arctan x-\sum_{k=0}^{n-1}(-1)^{k}\frac{x^{2k+1}}{2k+1}\Biggr],&x\neq 0;\\ 1,&x=0,\end{dcases}$
@@
$T_{2n}[\operatorname{arctanh}x]=\begin{dcases}\frac{2n+1}{x^{2n+1}}\Biggl(\operatorname{arctanh}x-\sum_{k=0}^{n-1}\frac{x^{2k+1}}{2k+1}\Biggr),&x\neq 0;\\ 1,&x=0\end{dcases}$
for $n\in\mathbb{N}_{0}$ and $x\in(-1,1)$. Among other findings, we derived the following conclusions:
  1. 1)
    For $n\in\mathbb{N}_{0}$, the normalized remainder $T_{2n}[\arcsin x]$ is a logarithmically absolutely monotonic function on $(0,1)$ and a logarithmically completely monotonic function on (1,0)(-1,0).
  2. 2)
    For $n\in\mathbb{N}_{0}$, the logarithmic function $\ln T_{2n}[\arcsin x]$ is absolutely monotonic on $(0,1)$ and completely monotonic on (1,0)(-1,0).
  3. 3)
    For $n\in\mathbb{N}_{0}$, the ratio $\frac{T_{2n+2}[\arcsin x]}{T_{2n}[\arcsin x]}$ is an absolutely monotonic function on $(0,1)$ and a completely monotonic function on (1,0)(-1,0).
  4. 4)
    For $n\in\mathbb{N}_{0}$, the normalized remainder $T_{2n}[\arctan x]$ is decreasing in $x\in(0,1)$.
  5. 5)
    For $n\in\mathbb{N}_{0}$, the normalized remainder T_2n[arctan(xi)]=T_2n[arctanhx] is a logarithmically absolutely monotonic function on $(0,1)$ and a logarithmically completely monotonic function on (1,0)(-1,0).
  6. 6)
    For $n\in\mathbb{N}_{0}$, the logarithmic function lnT_2n[arctan(xi)]=lnT_2n[arctanhx] is absolutely monotonic on $(0,1)$ and completely monotonic on (1,0)(-1,0), where $\operatorname{i}=\sqrt{-1}\,$ is the imaginary unit.
  7. 7)
    Let $x_{0}\in\bigl(-\frac{1}{2},-\frac{2}{5}\bigr)$ be the unique root of the equation
    $\sqrt{1-x^{2}}\,+x\arccos x=0,\quad x\in(-1,1).$ (3.15)
    Then the function $\arccos x$ is logarithmically concave on $(-1,x_{0})$ and logarithmically convex on $(x_{0},1)$.
  8. 8)
    The logarithmic functions
    $\ln\biggl(\frac{2}{\pi}\arccos x\biggr),\quad\ln T_{2n}[\arctan x],\quad\ln T_{2n}[\arcsin x]$
    were expanded into the Maclaurin power series.
@@
In [71, p. 71, Eq. (3.38)], we find the integral representation
@@
where the classical Euler gamma function $\Gamma(z)$ can be defined [71, Chapter 3] by
@@
This can also be derived by the Laurent series expansion of $\cot\theta$ in the handbooks [1, p. 75, Entry 4.3.70] and [44, p. 117, Entry 4.19.6]; see also [82, Remark 1]. As a result, the function $1-\frac{\theta}{\tan\theta}$ is absolutely monotonic on $(0,\pi)$ and is completely monotonic on $(-\pi,0)$.
@@
  1. 1)
    The normalized remainder $T_{2n+1}[\Phi(\theta)]$ for $n\in\mathbb{N}_{0}$ is positive in $\theta\in(-\pi,\pi)$, decreasing in $\theta\in(-\pi,0)$, and increasing in $\theta\in(0,\pi)$.
  2. 2)
@@
  3. 3)
    The normalized remainder $T_{2n+1}[\Phi(\theta)]$ for $n\in\mathbb{N}_{0}$ is an absolutely monotonic function in $\theta\in(0,\pi)$ and a completely monotonic function in $\theta\in(-\pi,0)$.
  4. 4)
@@
  5. 5)
    The ratio $\frac{T_{2n+3}[\Phi(\theta)]}{T_{2n+1}[\Phi(\theta)]}$ for $n\in\mathbb{N}_{0}$ is decreasing in $\theta\in(-\pi,0)$ and increasing in $\theta\in(0,\pi)$.
@@
    1. 1)
      is an absolutely monotonic function in $\theta\in(0,\pi)$ and a completely monotonic function in $\theta\in(-\pi,0)$,
    2. 2)
@@
The absolute and complete monotonicity of the ratio $\frac{T_{2n+3}[\Phi(\theta)]}{T_{2n+1}[\Phi(\theta)]}$ discovered in [68, Theorem 2] and the Laurent series of the function $\frac{1}{\Phi(\theta)}$ for $0<|\theta|<\pi$ expanded in [68, Theorem 3] imply that the function $\frac{2}{\theta^{2}}-\frac{1}{\Phi(\theta)}$ is absolutely monotonic on $(0,\pi)$ and completely monotonic on $(-\pi,0)$.
@@ 3.6. Normalized remainders of complete elliptic integrals
For given $x\in[0,1]$, the complete elliptic integral of the second kind $\mathscr{E}(x)$ are defined [71, p. 315] by
@@
In the paper [74, Remarks 2 and 4], we proved that the normalized remainders $T_{n}[\mathcal{H}(x)]$ and $T_{n}[\mathcal{G}(x)]$ for $n\in\mathbb{N}_{0}$ are absolutely monotonic on $(0,1)$.
@@
In the paper [24, Theorems 3 and 4], we employed a new method to concisely prove the absolute monotonicity of the normalized remainders $T_{n}[\mathcal{H}(x)]$ and $T_{n}[\mathcal{G}(x)]$ for $n\in\mathbb{N}_{0}$ on $(0,1)$ once again.
@@ 3.7. Normalized remainder of power of inverse sine function
In [17, Theorem 1], [18, Theorem 2.1], [52, Section 4], and [54, Section 6], we discovered the Maclaurin power series expansion
@@
for $n,\ell\in\mathbb{N}$ and the Stirling numbers of the first kind $s(\ell+n,n)$ for $\ell,n\in\mathbb{N}_{0}$ can be analytically generated [39, Theorem 3.14] by
@@
for $|x|<1$. These series expansions can be found in the literature [3, p. 444], [17, Remark 6], [18, Remark 5.2], [53, Section 1.1], [59, Section 1], [63, Section 1], [72, pp. 279 and 331–333], [85, p. 229], and [4, 8, 69], for example.
In [52, Theorem 7], see also [62, Example 3.1], the Maclaurin series expansion (3.17) was generalized as
@@
    where the central factorial numbers of the first kind $t(n,\ell)$ are defined [66, pp. 212–217, Section 6.5] by
@@ 4. Historic backgrounds in combinatorics and number theory
In this section, we collected some concepts, which can be reformulated in terms of normalized remainders, in combinatorics and number theory.
@@ 4.1. Generating functions of Carlitz–Howard’s generalizations of Bernoulli numbers and polynomials
The generalized Bernoulli polynomials $B_{n}^{(\sigma)}(t)$ can be analytically generated [71, p. 4] by
@@
for $n\in\mathbb{N}$, and they examined a lot of algebraic properties of these sequences and polynomials. It is easy to see that $A_{k}(0)=A_{k}$, $A_{1,k}=B_{k}$, $A_{2,k}=A_{k}$, $A_{n,k}(0)=A_{n,k}$, $A_{1,k}(0)=B_{k}$, $A_{2,k}(0)=A_{k}$, and $A_{1,k}(t)=B_{k}(t)$ for $k\in\mathbb{N}_{0}$. The sequence $A_{k}$ for $k\in\mathbb{N}_{0}$ was recently investigated in [5, Remark 2] and [64, Section 2]. In terms of the notation $T_{n-1}[\operatorname{e}^{x}]$, the equation (4.3) can be reformulated as
@@
See also [21, Remark 4], [22, Remark 4], [33, pp. 1036–1037], and [74, Remark 2].